Given 35.9million as working age population, 4.6million are part time workers and 15.71million working full time. 3.2million job search happened in less than 2 weeks ago among non workers, 1.72million most recently looked for work between 2 & 4 weeks ago. An additional 0.86million most recently looked for work 5weeks ago and 9.81m who don't have jobs have not looked for work in the past 6 weeks.

1. Size of the total labour force

2.How many people are unemployed?

3. labour force participation rate?

4. What is the unemployment rate?
1
Expert's answer
2020-04-01T08:52:38-0400

1.Total labor force is 4.6+15.71+3.2+1.72= 25.23 million people(those, who are working and looking for a job in last 4 weeks).

2.There are 3.2+1.72= 4.92 million people unemployed(those, who are not employed, but are activly looking for a job).

3.Labor force participation rate is "employed(and\ looking\ for\ job\ in\ past\ 4\ weeks)/total\ labor\ force" = 25.23/35.9= 70.27%

4.Unemloyment rate is "unemployed/civil\ labour\ force" =(3.2+1.72)/35.9= 13.7%